A car travels 540km in 4.5 hours.how far will it go in 8 hours at the same average speed?


1
Expert's answer
2021-05-30T13:20:53-0400

"V= \frac{L}{T} \\

V= \frac{540}{4.5}=120 \;km/hour \\

L = V \times T \\

L= 120 \times 8 = 960 \;km"

Answer: 960 km